OpenStudy (anonymous):

how to solve -x+6-5x=14-2x Steps

OpenStudy (anonymous):

So, first, you combine like terms: -6x + 6 = 14 - 2x Then move all the x's onto one side: 6 = 14 + 4x Then move the constants onto the other side: -8 = 4x Then isolate x: x = -2 Let me know if any of that is unclear.